The amount of an investment of P dollars for t years at simple interest rate of r is given by A = P + Prt. Rewrite this formular by factoring out the GCF on the right hand side. Also, find A if $8300 is invested for 3 years at a simple interest rate of 15%.

Answer by stanbon(75887)

A = P + Prt. Rewrite this formular by factoring out the GCF on the right hand side. Also, find A if $8300 is invested for 3 years at a simple interest rate of 15%.
---------
A=P(1+rt)
A=8300(1+0.15*3)
A=8300(1.45)
A=$12,035
Cheers,
Stan H.